OpenStudy (gschibby):

Can anyone help me with this? Solve this equation on y: (2x-3x^2)/2y = 0

OpenStudy (gschibby):

But doesn't that mean that when you multiply both sides with 2y (2y*0) that side becomes zero? Is that "legal"? :P

jigglypuff314 (jigglypuff314):

whoops, sorry, then I'm not sure how to do this either... sorry :/

OpenStudy (gschibby):

Ah well, thanks anyway :)

OpenStudy (gschibby):

I've got another example: \[-2x-y \div x+2y = 0\] and this is supposed to come out as y = -2x but how?

OpenStudy (anonymous):

\[\frac{2x-3x^{2}}{2y} = \frac{x(2-3x)}{2y}=0\] Your only y solution is that y ≠ 0 :P Your x solutions are that x = 0, and x = 2/3

OpenStudy (gschibby):

Oh, crap! I meant x, of course! But how do you get 2/3?

jigglypuff314 (jigglypuff314):

factor out the x in the numerator then the x in the numerator and the x in the denominator cancel out

OpenStudy (anonymous):

It's like a normal quadratic, just hiding :) you have x(2-3x) = 0, so x = 0 and (2 - 3x) =0

OpenStudy (gschibby):

Great! Thanks!^^

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

There isn't an x in the denominator ... :/ ?

OpenStudy (anonymous):

When you said, "Oh! I mean x" Gschibby you meant that you wanted the solutions for x right? Not that the 2y was supposed to be 2x? Or I am teh crzy?

OpenStudy (gschibby):

No no, I meant I wanted the solutions for x ;)

OpenStudy (gschibby):

You're absolutely right!^^

OpenStudy (anonymous):

^^

OpenStudy (gschibby):

I just posted another example, wich is the one I think jigglypuff answered ;)

OpenStudy (anonymous):

gotchya

OpenStudy (anonymous):

btw, those two solutions aren't like normal quadratic solutions. it's not a really a parabola....